Water drops fall at regular intervals from a top which is 5 m above the ground. The third drop is leaving the tap at the instant the first drop touches the ground. How far above the ground is the second drop at that instant.

Grade:12th pass

1 Answers

Venu Choudhary
5 years ago
First u need to find the time of flight of 1st drop as:T=sqrt(2h/g)=1sec
The time 1sec is divided in intervals for 3 drops as T/3..
So when the 1st drop has travelled for T/3 seconds then the second drop leaves..
Time of flight of 2nd drop when 1st drop reaches ground =T-T/3
Height above ground=5-1/2*g*t*t where t= T-T/3